2. A sales woman earns a fixed salary of $8000 a month plus commission of 5% on whatever amount she sells over $2500 in merchandise. What did she earn the month she sold $11,000 in merchandise?
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(57977)

A sales woman earns a fixed salary of $8000 a month plus commission of 5% on whatever amount she sells over $2500 in merchandise. What did she earn the month she sold $11,000 in merchandise?
---------------------------
E = 8000+0.05*(11000-2500)
----
E = 8000 + 0.05*8500
---
E = 8000 + 425
---
E = $1225
============
cheers,
Stan H.

Answer by ewatrrr(10682)

 
Hi
fixed salary of $8000 a month, plus 5% commission (over $2500 in sales)
What did she earn the month she sold $11,000 in merchandise?
$8000 + .05*$8500 = $8000 + $425 = $8425.